### v1

@@ 5.1 European/North Sea surges of January and February 2020
January and February 2020 was a particularly active storm period in Northern Europe with several large storm surges. A series of powerful low pressure weather systems passed through Northern Europe from west to east, creating prolonged periods of high water level along the coast, in particular around the North Sea basin. In Norway, two storms during this time period were classified as “extreme weather events” and consequently were given names. These were the storms Didrik on 15 January 2020 (Skjerdal et al., 2020), and Elsa 10–11 February 2020 (Selberg et al., 2020). The storm Elsa was named Ciara in the United Kingdom (UK), see Jardine et al. 2023, but we choose here to refer to it as Elsa.
@@ 5.1.1 Storm Didrik
The storm Didrik hit Southern Norway on 15 January 2020 (Skjerdal et al., 2020) with a forecast warning of high water level along the coast. The low pressure system, with a central pressure as low as $945\,\mathrm{hPa}$, formed off the southern tip of Greenland on 12 January and moved eastward between Iceland and the UK the following days. It brought with it strong westerly to south-westerly winds in the northern North Sea, which in addition to the inverse barometer effect itself resulted in a large transport of water towards the western Norwegian coast. The maximum total water level observed was among the top 10 highest for the stations from Stavanger in the south west to Ålesund on the north-western coast (Kristensen et al., 2024).
@@
Figure 7: Heatmap scatterplot of the observed and modeled storm surge for the positions shown in the maps in Figure 6.
@@ 5.1.2 Storm Elsa
The storm Elsa during 10–11 February 2020 (Selberg et al., 2020; Jardine et al., 2023) was the remnants of a long-lived extratropical cyclone that re-intensified as it passed south of the southern tip of Greenland. The passage eastward followed much the same path as the storm Didrik a month earlier. The maps in Figure 9 show the average RMSE for the daily maximum storm surge for all stations around the North Sea for the time period 9–12 February. The difference between the noIB and diaIB experiments are only minor, while the dynIB experiment stand out from the two with slightly lower RMSE for the Skagerrak, Norwegian and offshore North Sea stations. The heatmap scatterplot in Figure 10 also point to the dynIB as the best performing experiment for the Elsa storm, with both higher correlation and smaller departures from the diagonal. The time series for the four selected stations in Figure 11 show how the dynIB, most often, does a better job than noIB and diaIB at simulating both the daily maximum and the hourly water level.
@@
Figure 10: Heatmap scatterplot of the observed and modeled storm surge for the positions shown in the maps in Figure 9.
@@ 5.2.1 Hurricane Michael
The category 5 Hurricane Michael, on its track from south to north through the Gulf of Mexico, hit the south coast of western Florida near Mexico Beach on the 10 October 2018 (II et al., 2019; Kennedy et al., 2020). With winds up to $70\,\mathrm{m\,s^{-1}}$ and a central pressure down to $919\,\mathrm{hPa}$, Hurricane Michael also brought with it an estimated $3\,\mathrm{m}$ storm surge (Kennedy et al., 2020).
Figure 12 shows the average RMSE for the daily maximum storm surge for the available records along the western Florida coast for our three water level parameters. It is evident how important the IB effect is, but there are no distinguishable differences between the diaIB and dynIB. The same is evident in the heatmap scatterplots in Figure 13, where the majority of the data points fall very close to the diagonal for the diaIB and dynIB, while the noIB has slightly larger departures from the diagonal. This is also reflected in the linear correlations, $0.89$ for the diaIB and dynIB and $0.84$ for noIB. The highest observed storm surge in our dataset is approximately $2.0\,\mathrm{m}$ above mean sea level (for the chosen time period), while the model only has a storm surge of approximately $1.25\,\mathrm{m}$, as seen in Figure 13 and the time series in Figure 14a. However, all the four selected stations in Figure 14 show a marked improvement of both the hourly, and the daily maximum storm surge, with the introduction of the IB effect, but with no clearly visible difference between diaIB and dynIB. We note that even if there is a definite improvement in the results and reduction in error by introducing the IB effect, the highest maximum water level is still significantly underestimated. At the Apalachicola station (see Figure 14a, the maximum water level in the model is approximately 1 meter below the observed maximum water level. This is likely caused by the known fact that the ERA5 dataset exhibit generally too low wind speed and not low enough core pressure for tropical cyclones (Dulac et al., 2024). Similar underestimation is likely to be expected if examining other tropical cyclone cases.
@@
Figure 13: Heatmap scatterplot of the observed and modeled storm surge for the positions shown in the maps in Figure 12 during Hurricane Michael, 7-11 October 2018.
@@ 5.2.2 Typhoon Mangkhut
Typhoon Mangkhut (Ying et al., 2022; Yang et al., 2019) formed over the western Pacific in the beginning of September 2018, and intensified on its path westward towards the northern Philippines, where it made landfall as a category 5 typhoon on 14 September. Moving further westward, the typhoon weakened to a category 1 typhoon before making landfall in southern China, just south of Hong Kong, on 16 September 2018.
@@
Figure 16: Heatmap scatterplot of the observed and modeled storm surge for the positions shown in the maps in Figure 15 for Typhoon Mankhut, 13-17 September 2018.
@@ 6 Operational use and implications for storm surge forecasting
For operational medium-range forecasts, the IFS is run as an ensemble prediction system (ENS, e.g. ECMWF, 2024a), comprising 50 perturbed forecasts and one unperturbed (‘control’) forecast. The perturbed forecasts are designed to represent uncertainties due to the initial conditions and the model integrations (‘model uncertainty’), thus providing a description of forecast probabilities.
The perturbed initial conditions are generated via an ensemble of 4D-Var data assimilations (EDA Buizza et al., 2008; Isaksen et al., 2010; Lang et al., 2019) and singular vectors (Leutbecher and Palmer, 2008), which are both added to a central 4D-Var analysis (Rabier et al., 2000). In addition, uncertainties in the ocean initial conditions are represented with the use of 11 perturbed ocean analyses (Zuo et al., 2024).
Model uncertainties are represented via the Stochastically Perturbed Parameterisations scheme (SPP: Leutbecher et al., 2024; Lang et al., 2021; Ollinaho et al., 2017), which introduces random perturbations to uncertain processes in the atmospheric physics parametrisations throughout the forecast integration.
To explore the impact of including the IB effect, we ran an ensemble experiment using the operational configuration of the IFS (experiment id io66): one unperturbed and 50 perturbed forecasts, to 15 days. The atmospheric model uses the cubic octahedral reduced gaussian grid TCo1279 (approximately 9km grid-spacing) with 137 vertical levels. The atmospheric model is coupled to the ECMWF Wave Model (e.g. ECMWF, 2024b), the ORCA025_Z75 configuration of NEMO V4.0 (as previously described), and the LIM2 sea-ice model (Fichefet and Maqueda, 1997).
@@ 6.1 Ensemble simulations of the North Sea storm Xaver
Storm Xaver moved slowly from west to east across Northern Europe during 5-7 December 2013 (Deutschländer et al., 2013; Kristensen et al., 2024). The strong northerly winds over the North Sea resulted initially in a convergence of water along the eastern coast of the United Kingdom due to Ekman transport. When the wind direction started to back left, this water started to move anti-clockwise along the North Sea coastline as a free Kelvin wave. The combined Kelvin wave with local amplifications due to Ekman transport by the wind, and the inverse barometer effect resulted in a storm surge among the top five highest recordings over the last 100 years in the German Bight according to Deutschländer et al. 2013.
@@
As the low pressure system moved further east, and the wind direction continued backing further left, the storm surge signal moved further through the North Sea up along the western coast of Denmark and into the Skagerrak area as a Kelvin wave (see Kristensen et al. 2024).
@@
The forecasts are initialized daily at 00 UTC from the operational IFS initial conditions for the period from 2013-11-23 to 2013-12-07. To capture the IB effect, the experiment uses a research branch based on IFS CY49R2. The IB effect is estimated diagnostically from the mean sea level pressure from the IFS. We show how the coupled IFS ENS is able to forecast and give warning about the possibility of this very rare record event up to more than a week in advance in Figure 18. We selected two stations in the area of some of the highest observed storm surge during this event, namely the German station Cuxhaven, located in the German Bight in the southeastern North Sea, and the Danish station Hvide Sande, located at the western coast of Jutland, Denmark. The ensemble forecast provides an estimate of the uncertainty in a given forecast parameter via the ensemble spread: where the ensemble spread is largest, the uncertainty can be assumed to be greatest (and vice versa). In Figure 18, the ensemble spread is indicated by the shaded area that describes the ensemble range — the difference between the maximum value across the ensemble members and the minimum value. It is evident that by utilizing the entire ensemble, as opposed to only the control member (indicated by the solid lines with the round dots), there is a signal of a maximum surge level as large as the observation more than 5 days in advance at Cuxhaven and throughout the whole forecast period for Hvide Sande.
@@
The exceedance probability, defined as the number of ensemble members, as a percentage, exceeding a certain threshold (in this case, we used the maximum observed storm surge) for a given forecast lead time, is shown as the lower solid lines with square dots in both panels in Figure 18. For Cuxhaven, when we include the IB effect, the exceedance probability is larger than zero five days ahead of the event, and increases from three days ahead, to around $35\%$ by day zero. For Hvide Sande, with the IB effect included, the exceedance probability is above zero from more than 12 days ahead of the largest storm surge, with a strong increase from day five to above $50\%$ by day zero. By contrast, the forecasts that do not include the IB effect show exceedance probabilities at both locations of less than $10\%$ even at day zero.
@@
In operational forecasting applications, information from the real-time ENS runs can be assessed relative to the model climate, which is constructed from hindcasts (see e.g. Owens and Hewson, 2018, Section 5.3). This makes it possible to identify cases where the latest ENS forecasts are shifted towards extreme values. The stronger and earlier signal of large maximum storm surge levels from including the IB effect could ultimately provide even earlier warnings of extreme conditions than is indicated from our assessments of absolute forecast values.
@@
Figure 18: Probabilistic ensemble forecasts of surge levels for the 2013 North Sea Storm Xaver at the peak of the storm surge for two locations. Top panel (a) shows the German station Cuxhaven, located in the German Bight in the southeastern North Sea for maximum surge time (00UTC), while the bottom panel (b) shows the Danish station Hvide Sande, located at the western coast of Jutland, Denmark, for maximum surge time (06 UTC). The maximum observed storm surge during Storm Xaver for each station is indicated by the dotted red horizontal line. The blue and orange solid lines (with circular markers) show the control member with and without the inclusion of the IB effect, respectively. Similarly, the light blue and peach shaded areas show the ensemble range (from the lowest to the highest values from all ensemble members). The $y$-axis gives the storm surge relative to mean sea level [cm], while the $x$-axis indicates the forecast lead time (in days) prior to the time of the maximum observed storm surge. Please note that since there is only one daily run (starting at 00UTC) of the ENS, and the maximum storm surge occurs at different times at different stations, the forecast lead times will vary between the stations. In addition, the two lowest lines of each plot (with square markers) indicate the exceedance probability (with the maximum storm surge as criterion) in percent, as expressed by the right side $y$-axis, for the entire ensemble. The steel blue line is with the IB effect taken into account, while the chocolate brown line is without the IB effect.
@@ 7 Summary and conclusions
The purpose of the present work has been to evaluate two methods of including the inverse barometer effect in a stand-alone global ocean model. The model parameter under consideration is the weather-induced water level, and the two methods tested were to either add the IB effect diagnostically as a post process to a model run with constant sea level pressure (diaIB), or the direct dynamical inclusion during the model integration (dynIB).
@@
The modeled water level was compared against the Global Extreme Sea Level Analysis version 3 (GESLA3), a global dataset of water level observations. The general statistics that include all available model and observation data co-located in time and space yields a MAE of $11\,\mathrm{cm}$ for noIB, $10\,\mathrm{cm}$ for diaIB and $11\,\mathrm{cm}$ for dynIB. The corresponding RMSE values are $14\,\mathrm{cm}$, $13\,\mathrm{cm}$ and $14\,\mathrm{cm}$, respectively. The weather-induced contribution to the total water level is generally small. To better emphasize how well the three different parameters perform during larger storm surge events, we masked data where the absolute value of observed storm surge (relative to MSL) was less than one and two standard deviations. This obviously results in increased values for both MAE and RMSE for all three parameters (across all stations considered), with e.g., an RMSE increase of $4\,\mathrm{cm}$ for noIB, while both diaIB and dynIB increase by $2\,\mathrm{cm}$ for the masked data less that one standard deviation. When masking data where observations are less than two standard deviations, the RMSE increases with $11\,\mathrm{cm}$, $5\,\mathrm{cm}$ and $5\,\mathrm{cm}$ for the noIB, diaIB and dynIB, respectively, compared to the average RMSE over all measurements across all stations. In addition, the RMSE varies geographically between the stations and between the three parameters, as shown in the histograms in Figure 3. The geographical differences between the diaIB and dynIB are partly grouped, as seen in Figure 4 and 5 where the diaIB performs better than the dynIB on for example the European North West Shelf, and vice versa in the Baltic Sea. Generally, except for the Baltic Sea, the diaIB performs best, or the two perform equally. This is a key finding, since this demonstrates that all ocean models, even those that have no means of including the IB effect dynamically at integration time, can include the IB effect diagnostically as a post process to take into account the pressure effect on the storm surge water level.
@@
In addition to the general statistics, a series of case studies were analyzed to highlight the differences between the experiments for a few concrete cases. The time period covered by the model simulation contains several cases of strong extratropical and tropical cyclones. We analyze the two European winter storms Didrik and Elsa (the latter is known as storm Ciara in the UK), together with the cyclones Hurricane Michael and Typhoon Mangkhut. For the Didrik and Elsa cases, the combined statistics for all the selected stations indicate that the the dynIB is the best performing parameter with both highest correlation, and lowest RMSE. As seen in the maps in Figure 6 and 6, this is not the case for all stations, especially for the Skagerrak stations during Didirk, where the diaIB and noIB has lower RMSE than dynIB. Whereas the same stations exert lower RMSE for dynIB than diaIB and noIB during Elsa. Together with the previously mentioned geographical differences, this highlights how one of the alternative methods of including the IB effect is not always better than the other, and despite the fact that the diaIB yields the lowest average errors, local considerations should be taken into account. Based on the analysis of the two tropical cyclones, i.e. Hurricane Michael and Typhoon Mangkhut, there is a key point to be made, namely that although the inverse barometer effect must naturally be included to properly simulate and predict the change in water level due to storm surge, adding it diagnostically as a post process is a perfectly acceptable alternative for coarse models.
@@
To further assess and exemplify how the inclusion of the IB effect in a global ocean model can improve storm surge early warnings, we show the performance of storm surge derived from ECMWF ENS forecasts for Storm Xaver. In this case, the model was able to predict the possibility for the very rare and extreme event up to 12 days in advance when including the IB effect. We note that trying to accurately forecast and predict an extreme event such as Xaver is very challenging for a $0.25^{\circ}$ global model. Usually, the criterion for issuing warnings of high water level is linked to return periods, such as 5, 50 or 100 years (see e.g. Kristensen et al. 2022). So, the evaluation against the maximum observed storm surge for an event ranked top 5 over the last 100 years, as is done in the present work, should be considered a very strict criterion. The Storm Xaver case, and the use of the ensemble prediction system, highlights how the inclusion of the IB effect in the ENS forecasts with the global NEMO model makes the forecasts more capable of providing early warning of extreme storm surge events.
@@
To recapitulate, it is evident that NEMO, a global quarter degree resolution ocean model, forced with ERA5 winds and pressure fields is capable of reproducing quite faithfully the water level variations associated with the passage of both extratropical and tropical storms, even when compared against water level measurements (detided) in narrow fjords and estuaries. This is encouraging as it suggests that the current setup of ECMWF’s IFS can be used as boundary conditions for regional water level forecast systems with the simple addition of the inverse barometer effect. It also shows that an ensemble forecast system such as IFS can be used for early warnings of dangerous storm surges associated with tropical cyclones and extratropical storms.
@@
Observations of water level is obtained from the GESLA3 dataset (https://gesla787883612.wordpress.com/) (Haigh et al., 2023). A dataset containing the extracted time series from the NEMO ocean model runs can be downloaded from https://zenodo.org/records/16894653 (Mogensen and Kristensen, 2025). Additionally, the complete model experiments are archived at the ECMWF MARS archive as expid i5k6 and i5k8. Code used in the preparation of this manuscript include the python packages pangeo-pytide (https://github.com/CNES/pangeo-pytide) for performing harmonic analyses, Matplotlib (Hunter, 2007) for plotting and Cartopy (Met Office, 2010 - 2015) for creating maps. IPython notebooks for creating all the statistics and plots in this manuscript can be found here: https://github.com/nilsmkMET/nemo_sl_eval

### v2

@@ 5.1 European/North Sea surges of January and February 2020
January and February 2020 was a particularly active storm period in Northern Europe with several large storm surges. A series of powerful low pressure weather systems passed through Northern Europe from west to east, creating prolonged periods of high water level along the coast, in particular around the North Sea basin. In Norway, two storms during this time period were classified as “extreme weather events” and consequently were given names. These were the storms Didrik on 15 January 2020 (skjerdal:etal:2020), and Elsa 10–11 February 2020 (selberg:etal:2020). The storm Elsa was named Ciara in the United Kingdom (UK), see jardine2023, but we choose here to refer to it as Elsa.
@@ 5.1.1 Storm Didrik
The storm Didrik hit Southern Norway on 15 January 2020 (skjerdal:etal:2020) with a forecast warning of high water level along the coast. The low pressure system, with a central pressure as low as $945\,\mathrm{hPa}$, formed off the southern tip of Greenland on 12 January and moved eastward between Iceland and the UK the following days. It brought with it strong westerly to south-westerly winds in the northern North Sea, which in addition to the inverse barometer effect itself resulted in a large transport of water towards the western Norwegian coast. The maximum total water level observed was among the top 10 highest for the stations from Stavanger in the south west to Ålesund on the north-western coast (kristensen2024).
@@
Figure 7: Heatmap scatterplot of the observed and modeled residual water levels for the positions shown in the maps in Figure 6.
@@ 5.1.2 Storm Elsa
The storm Elsa during 10–11 February 2020 (selberg:etal:2020; jardine2023) was the remnants of a long-lived extratropical cyclone that re-intensified as it passed south of the southern tip of Greenland. The passage eastward followed much the same path as the storm Didrik a month earlier. The maps in Figure 9 show the average RMSE for the daily maximum storm surge for all stations around the North Sea for the time period 9–12 February. The difference between the noIB and diaIB experiments are only minor, while the dynIB experiment stand out from the two with slightly lower RMSE for the Skagerrak, Norwegian and offshore North Sea stations. The heatmap scatterplot in Figure 10 also point to the dynIB as the best performing experiment for the Elsa storm, with both higher correlation and smaller departures from the diagonal. The time series for the four selected stations in Figure 11 show how the dynIB, most often, does a better job than noIB and diaIB at simulating both the daily maximum and the hourly water level.
@@
Figure 10: Heatmap scatterplot of the observed and modeled residual water levels for the positions shown in the maps in Figure 9.
@@ 5.2.1 Hurricane Michael
The category 5 Hurricane Michael, on its track from south to north through the Gulf of Mexico, hit the south coast of western Florida near Mexico Beach on the 10 October 2018 (beven2019; kennedy2020). With winds up to $70\,\mathrm{m\,s^{-1}}$ and a central pressure down to $919\,\mathrm{hPa}$, Hurricane Michael also brought with it an estimated $3\,\mathrm{m}$ storm surge (kennedy2020).
Figure 12 shows the average RMSE for the daily maximum storm surge for the available records along the western Florida coast for our three water level parameters. It is evident how important the IB effect is, but there are no distinguishable differences between the diaIB and dynIB. The same is evident in the heatmap scatterplots in Figure 13, where the majority of the data points fall very close to the diagonal for the diaIB and dynIB, while the noIB has slightly larger departures from the diagonal. This is also reflected in the linear correlations, $0.89$ for the diaIB and dynIB and $0.84$ for noIB. The highest observed residual water level in our dataset is approximately $2.0\,\mathrm{m}$ above mean sea level (for the chosen time period), while the model only has a water level of approximately $1.25\,\mathrm{m}$, as seen in Figure 13 and the time series in Figure 14a. However, all the four selected stations in Figure 14 show a marked improvement of both the hourly, and the daily maximum storm surge, with the introduction of the IB effect, but with no clearly visible difference between diaIB and dynIB. We note that even if there is a definite improvement in the results and reduction in error by introducing the IB effect, the highest maximum water level is still significantly underestimated. At the Apalachicola station (see Figure 14a, the maximum water level in the model is approximately 1 meter below the observed maximum water level. This is likely caused by the known fact that the ERA5 dataset exhibit generally too low wind speed and not low enough core pressure for tropical cyclones (dulac2024). Similar underestimation is likely to be expected if examining other tropical cyclone cases.
@@
Figure 13: Heatmap scatterplot of the observed and modeled residual water levels for the positions shown in the maps in Figure 12 during Hurricane Michael, 7-11 October 2018.
@@ 5.2.2 Typhoon Mangkhut
Typhoon Mangkhut (ying2022; yang2019) formed over the western Pacific in the beginning of September 2018, and intensified on its path westward towards the northern Philippines, where it made landfall as a category 5 typhoon on 14 September. Moving further westward, the typhoon weakened to a category 1 typhoon before making landfall in southern China, just south of Hong Kong, on 16 September 2018.
@@
Figure 16: Heatmap scatterplot of the observed and modeled residual water levels for the positions shown in the maps in Figure 15 for Typhoon Mankhut, 13-17 September 2018.
@@ 6 Operational use and implications for storm surge forecasting
For operational medium-range forecasts, the IFS is run as an ensemble prediction system (ENS, e.g. ens), comprising 50 perturbed forecasts and one unperturbed (‘control’) forecast. The perturbed forecasts are designed to represent uncertainties due to the initial conditions and the model integrations (‘model uncertainty’), thus providing a description of forecast probabilities.
The perturbed initial conditions are generated via an ensemble of 4D-Var data assimilations (EDA buizza2008; isaksen2010; lang2019) and singular vectors (leutbecher2008), which are both added to a central 4D-Var analysis (rabier2000). In addition, uncertainties in the ocean initial conditions are represented with the use of 11 perturbed ocean analyses (zuo24).
Model uncertainties are represented via the Stochastically Perturbed Parameterisations scheme (SPP: leutbecher2024; lang2021; ollinaho2017), which introduces random perturbations to uncertain processes in the atmospheric physics parametrisations throughout the forecast integration.
To explore the impact of including the IB effect, we ran an ensemble experiment using the operational configuration of the IFS (experiment id io66): one unperturbed and 50 perturbed forecasts, to 15 days. The atmospheric model uses the cubic octahedral reduced gaussian grid TCo1279 (approximately 9km grid-spacing) with 137 vertical levels. The atmospheric model is coupled to the ECMWF Wave Model (e.g. wam), the ORCA025_Z75 configuration of NEMO V4.0 (as previously described), and the LIM2 sea-ice model (fichefet1997).
@@ 6.1 Ensemble simulations of the North Sea storm Xaver
Storm Xaver moved slowly from west to east across Northern Europe during 5-7 December 2013 (deutschlander:etal:2013; kristensen2024). The strong northerly winds over the North Sea resulted initially in a convergence of water along the eastern coast of the United Kingdom due to Ekman transport. When the wind direction started to back left, this water started to move anti-clockwise along the North Sea coastline as a free Kelvin wave. The combined Kelvin wave with local amplifications due to Ekman transport by the wind, and the inverse barometer effect resulted in a storm surge among the top five highest recordings over the last 100 years in the German Bight according to deutschlander:etal:2013.
@@
As the low pressure system moved further east, and the wind direction continued backing further left, the storm surge signal moved further through the North Sea up along the western coast of Denmark and into the Skagerrak area as a Kelvin wave (see kristensen2024).
@@
The forecasts are initialized daily at 00 UTC from the operational IFS initial conditions for the period from 2013-11-23 to 2013-12-07. To capture the IB effect, the experiment used a research branch based on IFS CY49R2. The IB effect was estimated diagnostically from the mean sea level pressure from the IFS. Figure 18 shows how the coupled IFS ENS is able to forecast and give warning about the possibility of this very rare record event up to more than a week in advance. We selected two stations in the area of the highest observed residual water level during this event, namely the German station Cuxhaven, located in the German Bight in the southeastern North Sea, and the Danish station Hvide Sande, located at the western coast of Jutland, Denmark. The ensemble forecast provides an estimate of the uncertainty in a given forecast parameter via the ensemble spread: where the ensemble spread is largest, the uncertainty can be assumed to be greatest (and vice versa). In Figure 18, the ensemble spread is indicated by the shaded area that describes the ensemble range — the difference between the maximum value across the ensemble members and the minimum value. It is evident that by utilizing the entire ensemble, as opposed to only the control member (indicated by the solid lines with the round dots), there is a signal of a maximum surge level as large as the observation more than 5 days in advance at Cuxhaven and throughout the whole forecast period for Hvide Sande.
@@
The exceedance probability, defined as the number of ensemble members, as a percentage, exceeding a certain threshold (in this case, we used the maximum observed residual water level) for a given forecast lead time, is shown as the lower solid lines with square dots in both panels in Figure 18. For Cuxhaven, when we include the IB effect, the exceedance probability is larger than zero five days ahead of the event, and increases from three days ahead, to around $35\%$ by day zero. For Hvide Sande, with the IB effect included, the exceedance probability is above zero from more than 12 days ahead of the largest storm surge, with a strong increase from day five to above $50\%$ by day zero. By contrast, the forecasts that do not include the IB effect show exceedance probabilities at both locations of less than $10\%$ even at day zero.
@@
In operational forecasting applications, information from the real-time ENS runs can be assessed relative to the model climate, which is constructed from hindcasts (see e.g. fug, Section 5.3). This makes it possible to identify cases where the latest ENS forecasts are shifted towards extreme values. The stronger and earlier signal of large maximum storm surge levels from including the IB effect could ultimately provide even earlier warnings of extreme conditions than is indicated from our assessments of absolute forecast values.
@@
Figure 18: Probabilistic ensemble forecasts of residual water levels for the 2013 North Sea Storm Xaver at the peak of the storm surge for two locations. Top panel (a) shows the German station Cuxhaven, located in the German Bight in the southeastern North Sea for maximum surge time ($0$ UTC), while the bottom panel (b) shows the Danish station Hvide Sande, located at the western coast of Jutland, Denmark, for maximum surge time (06 UTC). The maximum observed residual water level during Storm Xaver for each station is indicated by the dotted red horizontal line. The blue and orange solid lines (with circular markers) show the control member with and without the inclusion of the IB effect, respectively. Similarly, the light blue and peach shaded areas show the ensemble range (from the lowest to the highest values from all ensemble members). The $y$ axis gives the residual water level relative to mean sea level [cm], while the $x$ axis indicates the forecast lead time (in days) prior to the time of the maximum observed residual water level. Please note that since there is only one daily run (starting at 00UTC) of the ENS, and the maximum storm surge occurs at different times at different stations, the forecast lead times will vary between the stations. In addition, the two lowest lines of each plot (with square markers) indicate the exceedance probability (with the maximum storm surge as criterion) in percent, as expressed by the right side $y$ axis, for the entire ensemble. The steel blue line is with the IB effect taken into account, while the chocolate brown line is without the IB effect.
@@ 7 Summary and conclusions
The purpose of the present work has been to evaluate two methods of including the inverse barometer effect in a stand-alone global ocean model and to assess the potential use of including the IB effect diagnostically in the operational IFS ENS. The model parameter under consideration is the weather-induced water level, and the two methods tested were to either add the IB effect diagnostically as a post process to a model run with constant sea level pressure (diaIB), or the direct dynamical inclusion during the model integration (dynIB).
@@
The modeled water level was compared against the Global Extreme Sea Level Analysis version 3 (GESLA3), a global dataset of water level observations. The general statistics that include all available model and observation data co-located in time and space yields a MAE of $11\,\mathrm{cm}$ for noIB, $10\,\mathrm{cm}$ for diaIB and $11\,\mathrm{cm}$ for dynIB. The corresponding RMSE values are $14\,\mathrm{cm}$, $13\,\mathrm{cm}$ and $14\,\mathrm{cm}$, respectively. The weather-induced contribution to the total water level is generally small. To better emphasize how well the three different parameters perform during storm surge events, we masked data where the absolute value of observed residual water level (relative to MSL) was less than one and two standard deviations. This obviously results in increased values for both MAE and RMSE for all three parameters (across all stations considered), with e.g., an RMSE increase of $4\,\mathrm{cm}$ for noIB, while both diaIB and dynIB increase by $2\,\mathrm{cm}$ for the masked data less that one standard deviation. When masking data where observations are less than two standard deviations, the RMSE increases with $11\,\mathrm{cm}$, $5\,\mathrm{cm}$ and $5\,\mathrm{cm}$ for the noIB, diaIB and dynIB, respectively, compared to the average RMSE over all measurements across all stations. In addition, the RMSE varies geographically between the stations and between the three parameters, as shown in the histograms in Figure 3. The geographical differences between the diaIB and dynIB are partly grouped, as seen in Figure 4 and 5 where the diaIB performs better than the dynIB on for example the European North West Shelf, and vice versa in the Baltic Sea. Generally, except for the Baltic Sea, the diaIB performs best, or the two perform equally. This is a key finding, since this demonstrates that all ocean models, even those that have no means of including the IB effect dynamically at integration time, can include the IB effect diagnostically as a post process to take into account the pressure effect on the atmospherically induced water level variations.
Our finding that the diagnostic IB addition is often as good as dynamical inclusion globally is consistent with the benchmarks set by ponte1991 and wunsch1997, who found that global departures from isostatic behavior are generally small ($1-3~\mathrm{cm}$). However, the superiority of our dynIB experiment in enclosed basins, and during specific storm events, reflects the non-isostatic responses, such as resonance and restricted flow through choking points, documented by carrere2003 and gomis2008.
@@
In addition to the general statistics, a series of case studies were analyzed to highlight the differences between the experiments for a few concrete cases. The time period covered by the model simulation contains several cases of strong extratropical and tropical cyclones. We analyze the two European winter storms Didrik and Elsa (the latter is known as storm Ciara in the UK), together with the cyclones Hurricane Michael and Typhoon Mangkhut. For the Didrik and Elsa cases, the combined statistics for all the selected stations indicate that the the dynIB is the best performing parameter with both highest correlation, and lowest RMSE. As seen in the maps in Figure 6 and 9, this is not the case for all stations, especially for the Skagerrak stations during Didirk, where the diaIB and noIB has lower RMSE than dynIB. Whereas the same stations exert lower RMSE for dynIB than diaIB and noIB during Elsa. Together with the previously mentioned geographical differences, this highlights how one of the alternative methods of including the IB effect is not always better than the other, and despite the fact that the diaIB yields the lowest average errors, local considerations should be taken into account. Based on the analysis of the two tropical cyclones, i.e., Hurricane Michael and Typhoon Mangkhut, there is a key point to be made, namely that although the inverse barometer effect must naturally be included to properly simulate and predict the change in water level due to storm surge, adding it diagnostically as a post process is a perfectly acceptable alternative for coarse models.
@@
To further assess and exemplify how the inclusion of the IB effect in a global ocean model can improve storm surge early warnings, we show the performance of water level predictions derived from ECMWF ENS forecasts for Storm Xaver. In this case, the model was able to predict the possibility for the very rare and extreme event up to 12 days in advance when including the IB effect. We note that trying to accurately forecast and predict an extreme event such as Xaver is very challenging for a $0.25^{\circ}$ global model. Usually, the criterion for issuing warnings of high water level is linked to return periods, such as 5, 50 or 100 years (see e.g. kristensen22). So, the evaluation against the maximum observed storm surge for an event ranked top 5 over the last 100 years, as is done in the present work, should be considered a very strict criterion. The Storm Xaver case, and the use of the ensemble prediction system, highlights how the inclusion of the IB effect in the ENS forecasts with the global NEMO model makes the forecasts more capable of providing early warning of extreme storm surge events.
@@
To recapitulate, it is evident that NEMO, a global quarter degree resolution ocean model, forced with ERA5 winds and pressure fields, is capable of reproducing quite faithfully the water level variations associated with the passage of both extratropical and tropical storms, even when compared against water level measurements (detided) in narrow fjords and estuaries. This is encouraging as it suggests that the current setup of ECMWF’s IFS can be used as boundary conditions for regional water level forecast systems with the simple addition of the inverse barometer effect. It also shows that an ensemble forecast system such as IFS can be used for early warnings of dangerous storm surges associated with tropical cyclones and extratropical storms.
@@
Declaration of use of generative AI
During the preparation of this work the authors used Google Gemini in order to research literature during the revision process. After using this service, the authors reviewed and edited the content as needed and takes full responsibility for the content of the published article.
@@
Observations of water level is obtained from the GESLA3 dataset (https://gesla787883612.wordpress.com/) (haigh2023). A dataset containing the extracted time series from the NEMO ocean model runs can be downloaded from https://zenodo.org/records/16894653 (mogensen_2025). Additionally, the complete model experiments are archived at the ECMWF File System (ECFS) archive as expid i5k6 and i5k8. The ensemble experiment is archived as expid io66. Code used in the preparation of this manuscript include the python packages pangeo-pytide (https://github.com/CNES/pangeo-pytide) for performing harmonic analyses, Matplotlib (Matplotlib) for plotting and Cartopy (Cartopy) for creating maps. IPython notebooks for creating all the statistics and plots in this manuscript can be found here: https://github.com/nilsmkMET/nemo_sl_eval.